a gardener wishes to use 600 feet of fencing to enclosd a rectangular region and subdivide the region into two smaller rectangles . the total enclosed area is 15,000 square feet. find the dimensions of the enclosed region
Answer by TimothyLamb(4379)

---
2x + 2y = 600
xy = 15000
---
2x + 2y = 600
2y = 600 - 2x
y = 300 - x
---
xy = 15000
x(300 - x) = 15000
300x - xx = 15000
-xx + 300x - 15000 = 0
---
the above quadratic equation is in standard form, with a=-1, b=300, and c=-15000
---
to solve the quadratic equation, by using the quadratic formula, copy and paste this:
-1 300 -15000
into this solver: https://sooeet.com/math/quadratic-equation-solver.php
---
the two real roots (i.e. the two x-intercepts), of the quadratic are:
x = 63.3974596
x = 236.60254
---
answer:
w= 63.3975 ft
L= 236.6025 ft
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php